The sum of three numbers is 69 . The second number is 9 less than the first . The third number is 2 times the second. What are the numbers?

First number=?
Second number=?
Third=?

Answer by edjones(8007)

Let f=1st, s=2nd, t=3rd
f+s+t=69
.
s=f-9
f=s+9
.
t=2s
.
s+9+s+2s=69
4s=60
s=15
f=24
t=30
.
Ed